OpenStudy (anonymous):

An apartment building contains 12 units consisting of one- and two-bedroom apartments that rent for $360 and $450 per month, respectively. When all units are rented, the total monthly rental is $4,950. What is the number of twobedroom apartments? A. 3 B. 4 C. 5 D. 6 E. 7

OpenStudy (anonymous):

@Buttercup16

OpenStudy (anonymous):

450x+360(12-x)=4950 450x +4320-360x =4950 90x+4320=4950 90x=630 x= 630/90 x=7 two bed room apartments

OpenStudy (anonymous):

E.

OpenStudy (anonymous):

thanks